OpenStudy (mkmkasim):

∫(x^3)/(sqrt((x^2)+4)

OpenStudy (mkmkasim):

I know you have to use trig substitution

zepdrix (zepdrix):

\[\Large\bf\sf \frac{1}{2}\int\limits \frac{2x}{\sqrt{x^2+4}}\left(x^2\;dx\right)\]I would rewrite the expression like this and then do Integration by Parts. Oh trig sub? Is that what you're learning about right now? We can do that method instead.

OpenStudy (mkmkasim):

yeah we are learning trig sub

zepdrix (zepdrix):

Factor a 4 out of each term under the root,\[\Large\bf\sf \int\limits\limits \frac{x^3}{\sqrt{4\left[\left(x/2\right)^2+1\right]}}dx\]Then pull the the 4 out of the root,\[\Large\bf\sf \frac{1}{2}\int\limits\limits\limits \frac{x^3}{\sqrt{\left(\frac{x}{2}\right)^2+1}}dx\]This puts it into an easier form for us to set up a trig sub.

zepdrix (zepdrix):

So trig sub will allow us to get rid of the addition under the root, making it easier to simplify. We have (something)^2 + 1. Which trig relationship should we use? Do you remember which trig identity involves (something)^2 + 1 = (something else)^2

OpenStudy (mkmkasim):

1+tan^2 =sec^2

zepdrix (zepdrix):

Ok good good. We'll want to use tangent. That will allow us to simplify it down to secant^2 under the root. So we're making the substitution:\[\Large\bf\sf \frac{x}{2}=\tan \theta\]

zepdrix (zepdrix):

We'll need to replace the x in the numerator, and also the differential dx as well.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{1}{2}\int\limits\limits\limits\limits \frac{x^3}{\sqrt{\left(\color{orangered}{\frac{x}{2}}\right)^2+1}}dx\quad=\quad \frac{1}{2}\int\limits\limits\limits\limits \frac{x^3}{\sqrt{\left(\color{orangered}{\tan \theta}\right)^2+1}}dx\] \[\Large\bf\sf =\quad \frac{1}{2}\int\limits\limits\limits\limits \frac{x^3}{\color{orangered}{\sec \theta}}dx\]Ok so we've dealt with the denominator so far. That part make sense? :o

OpenStudy (mkmkasim):

yep so far so good

zepdrix (zepdrix):

So if this is our substitution, we can solve for x pretty easily:\[\Large\bf\sf \frac{x}{2}=\tan \theta\qquad\to\qquad x=2 \tan \theta\]What do you get for your differential?\[\Large\bf\sf dx=?\]

OpenStudy (mkmkasim):

2(secx)^2

zepdrix (zepdrix):

\[\Large\bf\sf dx=2\sec^2\theta\;d \theta\]Ok good. So now we just plug all the pieces in and simplify.

zepdrix (zepdrix):

\[\large\bf\sf =\quad \frac{1}{2}\int\limits\frac{(2\tan \theta)^3}{\sec \theta}\left(2\sec^2\theta\;d \theta\right)\quad = \quad 8\int\limits \frac{\tan^3 \theta}{\sec \theta}\left(\sec^2\theta\;d \theta\right)\]

zepdrix (zepdrix):

Hopefully I plugged those in correctly :o

OpenStudy (anonymous):

a u-sub will do nicely

OpenStudy (mkmkasim):

so then it becomes 8∫(tanx)^3 * secx

OpenStudy (mkmkasim):

so 8∫(tanx)^2(tanx)(secx)....then I got stuck

zepdrix (zepdrix):

Umm ya that one is a little tricky. I think one approach is to remember that: \(\Large\bf\sf (\sec x)'\quad=\quad \sec x\tan x\) We might be able to use this. Hmm

zepdrix (zepdrix):

Rewrite two of the tangent powers in terms of secant:\[\Large\bf\sf =8\int\limits \tan \theta (\sec^2\theta - 1)\sec \theta\;d \theta\]\[\Large\bf\sf =8\int\limits \sec^2\theta(\sec \theta \tan \theta)\;d \theta- 8\int\limits \sec \theta \tan \theta \;d \theta\]

OpenStudy (mkmkasim):

how did you go from the first part to the second?

zepdrix (zepdrix):

\[\Large\bf\sf =8\int\limits\limits \tan \theta \sec \theta(\sec^2\theta - 1)\;d \theta\]I distributed the tangent and secant to each term in the brackets. Then wrote it as two separate integrals.

OpenStudy (mkmkasim):

oh ok i see

zepdrix (zepdrix):

\[\Large\bf\sf =8\int\limits\limits \sec^2\theta\color{orangered}{(\sec \theta \tan \theta\;d \theta)}- 8\int\limits\limits \color{orangered}{(\sec \theta \tan \theta\;d \theta)}\]Understand how to solve those integrals? In each case, the orange part is the derivative of secant. So in the first one, you can do a u-sub to make it a bit clearer.

OpenStudy (mkmkasim):

so is is (8/3)(secx)^3 - 8secx

zepdrix (zepdrix):

Ya looks good. Then we need to get it back in terms of x, which can be a little tricky. We need to go back to our good ole triangle math.

zepdrix (zepdrix):